OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that passes through the points ( -2, 5) and (-2, 11). Do not use any spaces. Slope should be a reduced improper fraction.

OpenStudy (anonymous):

find slope between points (-2,5) and (-2,11) m = (y2 - y1) / (x2 - x1) where you have (x1, y1) and (x2, y2) for points after getting m, plug the value for m and any point into an equation to find the line y - y1 = m(x - x1) using (-2,5) for instance as (x1,y1) and the previously gotten value for m

OpenStudy (anonymous):

im confused